# Employing Sentence Space Embedding for Classification of Data Stream from Fake News Domain

### v1

@@ Abstract
Tabular data is considered the last unconquered castle of deep learning, yet the task of data stream classification is stated to be an equally important and demanding research area. Due to the temporal constraints, it is assumed that deep learning methods are not the optimal solution for application in this field. However, excluding the entire – and prevalent – group of methods seems rather rash given the progress that has been made in recent years in its development. For this reason, the following paper is the first to present an approach to natural language data stream classification using the sentence space method, which allows for encoding text into the form of a discrete digital signal. This allows the use of convolutional deep networks dedicated to image classification to solve the task of recognizing fake news based on text data. Based on the real-life Fakeddit dataset, the proposed approach was compared with state-of-the-art algorithms for data stream classification based on generalization ability and time complexity.
@@ 1 Introduction
There is a widespread opinion among researchers dealing with stream learning that using deep neural networks in this field is a suboptimal solution due to the data processing time [4]. For this reason – contrary to the opinion presented in the literature – tabular data seems not to be the only unconquered castle of deep learning [14]. The literature shows that neural networks are being gradually adapted to process data streams [10]. However, it remains one of the current research problems in the field of deep learning [4].
Excluding the entire pool of solutions characterized often by better quality than classical methods by assuming too high time complexity as the main argument may be hasty. Currently, networks are becoming increasingly popular, mainly thanks to tools such as Chat-GPT, DALL-e and Sora [32].
@@
The following work closely examines using deep networks in processing of data streams. The research was based on state-of-the-art methods dedicated to streaming data to ensure batch processing and take into account common problems affecting streams, such as the need for active model training or the occurrence of concept drift [17].
@@
This problem is juxtaposed with the second task, which is particularly important in the era of general availability of large language models, i.e., the recognition of fake news which are understood here as content constructed in a way that deliberately misleads the recipient to benefit the distributor. Online portals and social media platforms cause a flood of information, which requires stream processing, so models have a chance to adapt to new facts and language dynamics [19]. Machine learning models cannot process text data in their raw form, which poses an additional challenge to processing time in stream learning.
@@
In conjunction with neural networks for Natural Language Processing (nlp) tasks, embeddings are almost always used [13]. It is an advantage of embeddings over canonical methods based on n-grams, which maintain the semantic connection only with words in the local context. Nevertheless, there are methods based on non-neural network solutions in the literature that allow for achieving high quality in the classification of fake news [20].
@@
However, in the end, text preprocessing methods transform it into tabular data with a specific representation. There are models dedicated to text recognition problems that are adapted to such tasks and offer pre-trained weights [8]. However, the pool of such solutions is still smaller than the pool of methods available for image processing, for which convolutional networks are the primary classification tool [23]. The situation is similar with canonical tabular data, so methods that transform tables into images have already been proposed and used with deep networks have given promising results. This became the motivation for using this solution for text data in the following paper. For this purpose, the sentence space [15] representation was chosen, which allows the creation of an image even in the case of short texts, such as article titles.
Main contributions of this work are:
  • proposal of the Streaming Sentence Space (SSS) approach to – for the first time in the literature – use sentence space encoding to classify data streams containing text,
@@
  • developing the field of application of deep learning in data stream classification task, which is now considered one of the main research directions related to deep neural networks,
  • a comparison of SSS with state-of-the-art data stream ensemble classification algorithms in terms of classification accuracy and computational complexity.
@@ 2 Related works
This section presents the foundations of the proposed solution, both from the point of view of nlp and data stream processing.
@@ 2.1 Text data extraction methods
At the core of Natural Language Processing lies the challenge of converting natural language content into a numerical feature space that preserves the document’s semantic information. The bag-of-words method, a fundamental text feature extraction technique, calculates specific word occurrences within individual samples [21]. This approach forms the basis for more advanced methods such as bag-of-n-grams, which retain the contextual information of specific words and thereby a certain semantic association [11]. The tf-idf method [33] further enhances this vector-based approach by incorporating two key statistics: Term Frequency (tf), which measures word frequency within a document, and Inverse Document Frequency (idf), a logarithmic measure of word uniqueness within a corpus.
@@
In pursuit of the need to reduce dimensionality and the feature vector notation in continuous space, the continuous bag-of-words approach was invented, which, along with the continuous skip-gram model, is more widely known as Word2Vec (W2V[25]. This method allows the determination of a vector representation of a given length for each word in the corpus. Despite its innovation, this approach (i) fails to handle languages that are highly morphologically rich, and (ii) determines embeddings only based on local word relationships in the corpus. The answer to the first problem is the FastText approach [3], which, relying on the Word2Vec idea, performs additional word segmentation into character-n-grams, significantly enhancing the context during processing. In response to the second drawback, the Global Vectors (GloVe[31] model was developed, which is also based on Word2Vec, but in addition, extends the model to include general statistics from the processed corpus using a global word-word co-occurrence matrix.
@@
Among vectorization methods, Large Language Models (llms) are at the top-notch among the most commonly used approaches [13]. They use an artificial neural network structure called a transformer as their basis [37]. Their significant advantage is the employment of self-attention heads that powerfully extend the context, and this, in combination with the processing of massive linguistic resources, generates promising text representations [26]. One of the most popular large language models is Bidirectional Encoder Representations from Transformers (bert[8], which is designed to pre-train deep bidirectional representations from unlabeled text. As a result, it can capture the dependencies present throughout the text. Additionally, there are no words but subword units called WordPieces at the foundation of the bert model.
@@
However, despite the many advantages of large language models, they have a fair amount of complexity, affecting the time required to determine vectors. A MiniLM model [40] is an interesting approach where the authors train a reduced model using knowledge distillation from the bert, maintaining a quality similar to the original. The primary transformer idea draws inspiration from the sequence processing concept – the seq2seq model [36]. The expected input is a sequence of tokens forming a longer text. In contrast, a more intuitive solution would be to employ word embeddings in processing a single word into a vector space.
@@ 2.2 Multi-Dimensional Encoding of text data
Access to massive volumes of data and increased processing power promotes the usage of deep learning techniques. They are the foundation of multimodal data processing, which is primarily reliant on computer vision tasks, where deep methods frequently outperform canonical approaches [29]. Numerous scientific articles confirm that convolutional networks are successfully utilized for image, video, natural language [12], and audio classification (e.g., in spectrogram form) [34] in both unimodal and multimodal settings. Deep networks facilitate transfer learning, enabling models to apply previously learned information to the task at hand [42].
Although the term multi-dimensional encoding is mainly used for tabular data, the Sentence Space proposed by Kim [15] can be considered as its equivalent for text corpora. Using sentence space, text data is transformed into an image in which each row contains embeddings of individual words for each text sample. This approach is clearly dependent on the configuration of the convolutional neural network architecture used, as noted and studied in their work by Zhang and Wallace [41], analyzing possible configurations of one-layer cnns. In turn, Le et al. [22] analyzed the effect of the depth of convolutional neural networks on sentence space. Lately, an extension of the original concept of text encoding to a two-dimensional discrete digital signal was proposed by Soni et al. [35] in the form of TextConvoNet. This approach extracts the n-gram features within a sentence and captures the n-gram features between sentences in the input text data, resulting in a three-dimensional representation.
@@ 2.3 Classifier ensemble for imbalanced data stream
Despite over three decades of research and numerous techniques now accessible, classifying drifting imbalanced data streams remains one of the important machine learning topics. Methods designed for this type of data can work in online manner, where each instance is analyzed individually, or on batches of data, where the stream is processed in non-interlacing windows. This work focuses on batch processing, which, because to the larger training set, may provide improved classification quality in the current concept, but is accompanied with a delayed reaction due to having to wait for the next data batch to be available [1].
@@
Data imbalance is a prevalent problem in data streams, where the imbalance ratio can be static or dynamic [2]. Most real streams do not have a fixed imbalance ratio, and their properties might change over time [39]. As a result, data stream classification methods should achieve good classification quality regardless of class distribution, however most techniques built for imbalanced data streams produce unsatisfactory results when class sizes are similar. At the same time, algorithms designed with balanced data in mind, frequently have difficulty with the correct classification of data streams with skewed class distribution [6]. Methods designed for for dealing with imbalanced data are separated into two main groups: (i) data-level approaches and (ii) algorithm-level techniques [1]. The first group focuses on data preprocessing, canonically by employing oversampling or undersampling, to change its characteristics prior to classification attempt to alleviate the bias towards majority class, whereas algorithm-level approaches focus on modifying classification algorithms’ training phase.
@@
The most prevalent methods for imbalanced data stream classification employ classifier ensembles coupled with data preprocessing techniques. By assuring diversity, constantly updating the classifier pool, and combining the available models, it is feasible to increase the generalization ability and allow for dynamic adaptation to changes in the stream’s characteristics [5]. Among the established algorithms, we can distinguish Learn++.CDS (Concept Drift with Smote) and Learn++.NIE (Nonstationary and Imbalanced Environments) by Ditzler and Polikar [9]. Learn++.CDS extends the Learn++.NIE (Non-Stationary Environments) algorithm by employing the smote in attempt to balance the number of samples in each class, while Learn++.NIE utilizes a penalty constraint to balance classification accuracy on all classes, while also employing a bagging-based sub-ensemble. Wang et al. proposed the Oversampling Online Bagging (oob) and Undersampling Online Bagging (uob) algorithms dedicated for online data stream processing, extending Online Bagging by altering the Poisson distribution $\lambda$ parameter according to the current imbalance ratio [38]. Cano and Krawczyk developed Kappa Updated Ensemble (kue[6], which combines batch-based and online processing on feature subspaces. kue uses the Kappa statistic to dynamically weight and select base classifiers. The same authors introduced also the Robust Online Self-Adjusting Ensemble (rose) for non-stationary data stream classification [7]. This method trains online learners based on data views, ensuring a diverse classifier pool. It also employs drift detectors to respond quickly to changes in data distribution and proposes effective strategies for dealing with data imbalance. Wozniak et al. used built-in mechanisms (e.g., weighting and aging of classification models) to establish a self-updating classifier pool that can adjust its lineup in response to changes in imbalance ratio and concept drift. Klikowski and Woźniak trained one-class classifiers using clustered data [16], while Zyblewski et al. proposed to combine Dynamic Classifier Selection with data preprocessing techniques for imbalanced data stream classification [43].
@@ 3 Streaming Sentence Space
As observed in the introduction of this paper, solutions based on deep learning architectures are often overlooked in data stream classification tasks. This is due to concerns about increased computational and time complexity, in both the induction and inference process [24], despite the tremendous recent progress made in the area of deep learning. One promising solution is Sentence Space, a multi-dimensional encoding counterpart for text data. Despite numerous works confirming the performance of sentence space and its derivatives, no studies analyzing the use of this encoding in the task of classifying streams containing text data have been produced so far. In order to fill this niche, this paper proposes Streaming Sentence Space (sss), thus taking the first step toward analyzing the application of sentence space in the classification of dynamically imbalanced data streams from fake news domain.
The main assumption behind the presented research was to keep the time complexity as low as possible and to enable the use of SSS in real-life data stream classification tasks while maintaining the generalization potential inherent to convolutional neural networks. Accordingly, this work analyzes only a batch-based processing scenario, in which prediction and model training are performed on a window of predefined size. The need to wait for a single data chunk to fill up, depending on the dynamics of the data stream, can significantly reduce problems arising from possibly increased processing time.
The basis of sss is the conventional sentence space encoding [15], in which individual words are transformed into embeddings that represent consecutive lines of an image. Of course, it is also possible to use approaches such as TextConvoNet, but this depends on the characteristics of the data being analyzed. The decision to use sentence space in this case was related to the relative short length of the texts contained in the stream corpus (more in Section 4). As for the convolutional network, the decision was made to use the popular ResNet-18 architecture with the assumption of only one training epoch in each successive data chunks. The standard and commonly employed optimizer sgd with learning rate of 0.001 and momentum of 0.9 was used, batch size was set to 8. Cross-entropy loss was used as the loss function.
@@
One of the fundamental problems arising from sentence space is the approach to determining the dimensions of the images resulting from encoding. The width here is typically the length of the embedding vector, but selecting the height is no longer a trivial task, as it depends on the number of words contained in the text. Approaches based on padding or clipping texts to a specific length can be used here, but in the case of sss, it was decided to resize the heights of the images by bilinear interpolation. The first two subplots of Fig 2 show the experimental process of selecting image heights for the Fakeddit dataset analyzed. Text embeddings were obtained using the GloVe technique, which is a more recent alternative to Word2Vec usually used for this purpose and offers better recognition quality in many problems. After analyzing the distribution of the number of words in the corpus texts, it was found that almost all of them were in the range of up to 50 words in length. Due to this observation, 50x300 px was set as the initial dimensions of sentence space images after resize. In addition, this experiment was repeated for dimensions of 100x300 px and 200x300 px. The obtained values of balanced accuracy score, although very close, indicate the advantage of images with a height of 200 px, and therefore this is the value used in the experiments presented next.
In addition, due to the relatively unusual characteristics of the resulting images of sentence space encoding and the possibility of negative knowledge transfer, a short experiment was conducted to determine the validity of using the ResNet-18 architecture pre-trained on the ImageNet dataset. In the image height experiment, transfer learning was applied by default, as is the case in many research articles, but in this case it was decided to repeat the study using the Resnet model architecture learned from scratch. The results obtained, presented in the last subplot of Fig 2, indicate a minimal advantage of the pre-trained network when classifying images resulting from sentence space encoding.

### v2

@@ Abstract
Due to the complexity constraints, it is assumed that deep learning methods are not the optimal solution for application in data stream classification tasks. However, excluding the entire – and prevalent – group of methods seems rather rash given the progress that has been made in recent years in its development. For this reason, the following paper is the first to present an approach to natural language data stream classification using the sentence space method, which allows for encoding text into the form of a discrete digital signal. This allows the use of convolutional deep networks dedicated to image classification to solve the task of recognizing fake news based on text data. Based on the real-life Fakeddit dataset, the proposed approach was compared with state-of-the-art algorithms for data stream classification based on generalization ability and time complexity.
@@ 1 Introduction
A widespread opinion among stream learning researchers says that using deep neural networks in this field is a suboptimal solution due to the data processing time [borisov2022deep]. Excluding the entire pool of solutions characterized often by better quality than canonical methods by assuming too high time complexity as the main argument may be hasty. Thus, the gradual process of deep learning adaptation for data stream processing can be noticed [duda2020training] in literature, even though currently, it remains an open problem [borisov2022deep].
@@
The following work focuses on employing convolutional neural networks to process natural language data streams. The research was based on state-of-the-art batch-based methods dedicated to streaming data to account for common problems affecting streams, such as the need for dynamic model updating or the occurrence of concept drift [komorniczak2023complexity].
@@
Natural language processing is crucial in the era of general availability of Large Language Models (llms). Online portals and social media platforms cause a flood of information and empower the spread of fake news. Analyzing all this data requires stream processing, so models have a chance to adapt to new facts and language dynamics [ksieniewicz2020fake]. Yet machine learning models cannot process text data in their raw form, which poses an additional challenge to processing time in stream learning.
@@
In conjunction with neural networks for Natural Language Processing (nlp) tasks, embeddings are almost always used [INCITTI2023418]. It is an advantage of embeddings over canonical methods based on n-grams, which only maintain semantic connection with words in the local context. Nevertheless, in literature, some methods allow for achieving high quality in classifying fake news without using neural networks [ksieniewicz2023alphabet].
@@
However, in the end, text preprocessing methods transform it into tabular data. There are models dedicated to text recognition problems that are adapted to such tasks and offer pretrained weights [devlin2018bert]. However, the pool of such solutions is still smaller than the pool of methods available for image processing, for which convolutional networks are the primary classification tool [li2021survey]. Methods that transform tabular data into images have already been proposed and used with deep networks with promising results [zyblewski2024employing]. This became the motivation for using this solution for text data in the following paper. For this purpose, the sentence space [kim2014convolutional] representation was chosen, which allows the creation of an image even in the case of short texts, such as article titles.
@@
Main contributions of this work are: (a) proposal of the Streaming Sentence Space (sss) novelty approach, as an adaptation of sentence space encoding for text data stream classification, (b) development the field of application of deep learning in data stream classification task, which is now considered one of the main research directions related to deep neural networks, and (c) comparison of SSS with state-of-the-art data stream ensemble classification algorithms in terms of classification accuracy and computational complexity.
@@ 2 Related works
This section presents the foundations of the proposed solution, both from the point of view of nlp and data stream processing, and introduces the reference state-of-the-art ensemble classification algorithms dedicated to chunk-based stream processing.
@@ 2.1 Text data extraction methods
At the core of Natural Language Processing lies the challenge of converting natural language content into a numerical feature space that preserves the document’s semantic information. The bag-of-words method, a fundamental text feature extraction technique, calculates specific word occurrences within individual samples [LANG1995331]. This approach forms the basis for more advanced methods such as bag-of-n-grams, which retain the contextual information of specific words and thereby a certain semantic association [furnkranz1998study]. The tf-idf method [salton1987term] further enhances this vector-based approach by incorporating two key statistics: Term Frequency (tf), which measures word frequency within a document, and Inverse Document Frequency (idf), a logarithmic measure of word uniqueness within a corpus.
@@
In pursuit of the need to reduce dimensionality and the feature vector notation in continuous space, the continuous bag-of-words approach was invented, which, along with the continuous skip-gram model, is more widely known as Word2Vec (W2V[mikolov2013efficient]. This method allows the determination of a vector representation of a given length for each word in the corpus. Despite its innovation, this approach (a) fails to handle languages that are highly morphologically rich, and (b) determines embeddings only based on local word relationships in the corpus. The answer to the first problem is the FastText approach [10.1162/tacl_a_00051], which, relying on the Word2Vec idea, performs additional word segmentation into character-n-grams, significantly enhancing the context during processing. In response to the second drawback, the Global Vectors (GloVe[pennington2014glove] model was developed, which is also based on Word2Vec, but in addition, extends the model to include general statistics from the processed corpus using a global word-word co-occurrence matrix.
@@
Large Language Models are top-notch among the most commonly used vectorization methods [INCITTI2023418]. They implement a neural network structure called a transformer as their basis [NIPS2017_3f5ee243]. Their significant advantage is the employment of self-attention heads that powerfully extend the context, and this, in combination with the processing of massive linguistic resources, generates promising text representations [10.1145/3605943]. One of the most popular large language models is Bidirectional Encoder Representations from Transformers (bert[devlin2018bert], which is designed to pre-train deep bidirectional representations from unlabeled text. As a result, it can capture the dependencies present throughout the text. Additionally, there are no words but subword units called WordPieces at the foundation of the bert model.
@@
However, despite the many advantages of large language models, they have a fair amount of complexity, affecting the time required to determine vectors. A MiniLM model [NEURIPS2020_3f5ee243] is an interesting approach where the authors train a reduced model using knowledge distillation from the bert, maintaining a quality similar to the original.
@@ 2.2 Multi-Dimensional Encoding of text data
Access to massive volumes of data and increased processing power promotes the usage of deep learning techniques. They are the foundation of multimodal data processing, which is primarily reliant on computer vision tasks, where deep methods frequently outperform canonical approaches [o2020deep]. Numerous scientific articles confirm that convolutional networks are successfully utilized for image, video, natural language [gimenez2020semantic], and audio classification (e.g., in spectrogram form) [satt2017efficient] in both unimodal and multimodal settings. Deep networks also facilitate transfer learning, enabling models to apply previously learned information to the task at hand [zhuang2020comprehensive].
Although the term Multi-Dimensional Encoding (mde) is mainly used for tabular data, the Sentence Space proposed by Kim [kim2014convolutional] can be considered as its equivalent for text corpora. Using sentence space, text data is transformed into an image in which each row contains embeddings of individual words for each text sample. This approach is clearly dependent on the configuration of the convolutional neural network architecture used, as noted and studied in their work by Zhang and Wallace [zhang2015sensitivity], analyzing possible configurations of one-layer cnns. In turn, Le et al. [le2018convolutional] analyzed the effect of the depth of convolutional neural networks on sentence space. Lately, an extension of the original concept of text encoding to a two-dimensional discrete digital signal was proposed by Soni et al. [soni2023textconvonet] in the form of TextConvoNet. This approach extracts the n-gram features within a sentence and captures the n-gram features between sentences in the input text data, resulting in a three-dimensional representation.
@@ 2.3 Classifier ensemble for imbalanced data stream
Despite over three decades of research and availability of numerous methods, classifying drifting imbalanced data streams remains one of the important machine learning topics. Methods designed for this task can work in online manner, where each instance is analyzed individually, or on batches of data, where the stream is processed in non-interlacing windows. This work focuses on batch processing, which, because of the larger training set, may provide improved classification quality in the current concept, but is accompanied with a delayed reaction due to waiting for the next data batch [aguiar2023survey].
@@
Data imbalance is a prevalent problem in data streams, where the imbalance ratio can be static or dynamic [aminian2019study]. Most real streams do not have a fixed imbalance ratio, and their properties might change over time [wang2018systematic]. As a result, data stream classification methods should achieve good classification quality regardless of class distribution, however most techniques built for imbalanced data streams produce unsatisfactory results when class sizes are similar. At the same time, algorithms designed with balanced data in mind, frequently have difficulty with the correct classification of data streams with skewed class distribution [cano2020kappa]. Methods designed for for dealing with imbalanced data are separated into two main groups: (a) data-level approaches and (b) algorithm-level techniques [aguiar2023survey]. The first group focuses on data preprocessing, canonically by employing oversampling or undersampling, to change its characteristics prior to classification attempt to alleviate the bias towards majority class, whereas algorithm-level approaches focus on modifying classification algorithms’ training phase.
@@
The most prevalent methods for imbalanced data stream classification employ classifier ensembles coupled with data preprocessing techniques. By assuring diversity, constantly updating the classifier pool, and combining the available models, it is feasible to increase the generalization ability and allow for dynamic adaptation to changes in the stream’s characteristics [brzezinski2018ensemble]. Among the established algorithms, we can distinguish Concept Drift with Smote (Learn++.CDS) and Nonstationary and Imbalanced Environments (Learn++.NIE) by Ditzler and Polikar [Ditzler:2013]. Learn++.CDS extends the Learn++.NIE algorithm by employing the smote in attempt to balance the number of samples in each class, while Learn++.NIE utilizes a penalty constraint to balance classification accuracy on all classes, while also employing a bagging-based sub-ensemble. Wang et al. [Wang2015] proposed the Oversampling Online Bagging (oob) and Undersampling Online Bagging (uob) algorithms dedicated for online data stream processing, extending Online Bagging by altering the Poisson distribution $\lambda$ parameter according to the current imbalance ratio. Cano and Krawczyk [cano2020kappa] developed Kappa Updated Ensemble (kue), which combines batch-based and online processing on feature subspaces. kue uses the Kappa statistic to dynamically weight and select base classifiers. The same authors introduced also the Robust Online Self-Adjusting Ensemble (rose) for non-stationary data stream classification [cano2022rose]. This method trains online learners based on data views, ensuring a diverse classifier pool. It also employs drift detectors to respond quickly to changes in data distribution and proposes effective strategies for dealing with data imbalance. Woźniak et al. used built-in mechanisms (e.g., weighting and aging of classification models) to establish a self-updating classifier pool that can adjust its lineup in response to changes in imbalance ratio and concept drift [wozniak2023active]. Klikowski and Woźniak [klikowski2020employing] trained one-class classifiers using clustered data, while Zyblewski et al. [zyblewski2021preprocessed] proposed to combine Dynamic Classifier Selection with data preprocessing techniques for imbalanced data stream classification.
@@ 3 Streaming Sentence Space
Figure 1: The general scheme of the proposed SSS approach.
As observed in Section 1, solutions based on deep learning are often overlooked in data stream classification tasks. This is due to concerns about increased computational and time complexity, in both the induction and inference process [michalski1993inferential], despite the tremendous recent progress made in the area of deep learning. One promising solution is Sentence Space, a multi-dimensional encoding counterpart for text data. While attempts have already been made in the literature to use mde approaches for tabular data in data stream scenarios [zyblewski2024employing], despite numerous works confirming the performance of sentence space and its derivatives, no studies analyzing the use of this encoding in the task of classifying streams containing text data have been produced so far. In order to fill this niche, this paper proposes Streaming Sentence Space (sss), thus taking the first step toward analyzing the application of sentence space in the classification of dynamically imbalanced data streams from fake news domain.
The main assumption behind this work was to keep the time complexity as low as possible and to enable the use of sss in real-life data stream classification tasks while maintaining the generalization potential inherent to convolutional neural networks. Accordingly, this work analyzes only a batch-based processing scenario, in which prediction and model training are performed on a window of predefined size. The need to wait for a single data chunk to fill up, depending on the dynamics of the data stream, can significantly reduce problems arising from possibly increased processing time.
The basis of sss is the conventional sentence space encoding [kim2014convolutional], in which individual words are transformed into embeddings that represent consecutive lines of an image. Of course, it is also possible to use approaches such as TextConvoNet, but this depends on the characteristics of the data being analyzed. The decision to use sentence space in this case was related to the relative short length of the texts contained in the stream corpus (more in Section 4). As for the convolutional network, the decision was made to use the popular ResNet-18 [he2015deepresiduallearningimage] architecture with the assumption of only one training epoch in each successive data chunks. The standard and commonly employed optimizer sgd with learning rate of 0.001 and momentum of 0.9 was used, batch size was set to 8. Cross-entropy loss was used as the loss function.
@@
One of the fundamental problems arising from sentence space is the approach to determining the dimensions of the images resulting from encoding. The width is typically the length of the embedding vector, but selecting the height is not a trivial task, as it depends on the number of words contained in the text. Approaches based on padding or clipping texts to a specific length can be used, but in the case of sss, it was decided to resize the heights of the images by bilinear interpolation. The first two subplots of Figure 2 show the experimental process of selecting image heights for the fakeddit dataset. Text embeddings were obtained using the GloVe technique, which is a more recent alternative to Word2Vec usually used for this purpose and offers better recognition quality in many problems. After analyzing the distribution of the number of words in the corpus texts, it was found that almost all of them were in the range of up to 50 words in length. Due to this observation, 50x300 px was set as the initial dimensions of sentence space images after resize. In addition, this experiment was repeated for dimensions of 100x300 px and 200x300 px. The obtained values of balanced accuracy score, although very close, indicate the advantage of images with a height of 200 px, and therefore this is the value used in the experiments presented next.
In addition, due to the relatively unusual characteristics of the resulting images of sentence space encoding and the possibility of negative knowledge transfer, a short experiment was conducted to determine the validity of using the ResNet-18 architecture pretrained on the ImageNet dataset. In the image height experiment, transfer learning was applied by default, as is the case in many research articles, but in this case it was decided to repeat the study using the ResNet-18 model architecture learned from scratch. The results obtained, presented in the last subplot of Fig. 2, indicate a minimal advantage of the pretrained network when classifying images resulting from sentence space encoding.